Divide 73 into two parts whose product is 402
Answer by amarjeeth123(569)

Let the two parts be x and (73-x)
Product=402
x(73-x)=402
73x-x^2=402
x^2-73x+402=0
x^2-67x-6x+402=0
x(x-67)-6(x-67)=0
(x-6)(x-67)=0
x=67 or x=6.
The two parts are 6 and 67.